A 5.20g bullet moving at 672 m/s strikes a 700g wooden block atrest on a frictionless surface. The bullet emerges, travelingin the same direction with its speed reduced to 428 m/s.
a. What is the resulting speed of the block?
b. What is the speed of the bullet-block center of mass?
Answer #1
We have:
Mass of bullet mb = 5.20 g
Speed of bullet initial (Vi)b = 672 m/s
mass of wooden block mbl = 700g
speed of block initial (Vi)bl = 0
final speed of bullet (Vf)b = 428m/s

final speed of block (Vf)bl = ?
given by law of conservation of momentum
mb * (Vi)b + (mbl)*(Vi)bl = mb*(Vf)b + mbl *(Vf)bl
5.20g * 672m/s + 700g*0 = 5.20g*428m/s +700g * (Vf)bl
(Vf)bl = 1.81m/s

let theVcom be the speed of the bullet block center of massso,
Vcom = [mb* (Vi)b]/[mb + mbl]
          = [5.2 *672]/[5.2+700]
          = 4.955 m/s

Hope this helps
Answer #2
Actually, ^that^ is wrong. Everything is correct expect for thatthe final velocity of the bullet is 428 m/s, not (672-428) m/s. Itsays "reduced TO 428" not "reduced by 428"

So the final velocity of the block is acutally about 1.81 m/s
Answer #3
_   
      
a)
_   
=>     
=>   
=>      (m/s )